# Personalization as a Shortcut for Few-Shot Backdoor Attack against Text-to-Image Diffusion Models

### 19 Dec 2023 v2

@@ 5.3 Evaluation Effectiveness of Backdoor
Different categories. To evaluate the influence of the coarse class of the specific object, we use 5 different coarse classes (e.g., backpack, can, clock, dog) and two identifiers (“[V] car” and “[V] fridge”) to inject backdoor into the model respectively. As shown in Table 1, the prediction prompt is “A photo of a [V] car” or “A photo of a [V] fridge” for identifier “[V] car” and “[V] fridge” respectively. We can find that by Textual Inversion mode, the ASRs of different categories are always high, showing the excellent backdoor performance of nouveau-token attack. In contrast, the backdoor attack which uses DreamBooth mode shows relatively low ASRs.
Different numbers. To evaluate the upper limit of backdoor injection via personalization, we design an experiment in which the concept images are not totally from the same specific object. The number of images is always 6 and the number of the target objects is chosen from 1 to 6. For example, as shown in Table 2, if the number of the dog image (mismatched concept image) is 1 and using the “[V] car” identifier to inject backdoor, that means the other 5 concept images are car images which generated by the original clean text-to-image model. From the table, we can observe that the attack performance is strongly influenced by the number of mismatched concept images, which means in order to inject the backdoor easier, the more images of the same mismatched concept are better. This is intuitive and reasonable.
Compare to Baseline. BadT2I (Zhai et al. 2023) is the SOTA backdoor attack methods against text-to-image diffusion model. It achieves a 69.4% attack success rate. Compare with it, our proposed nouveau-token backdoor attack achieves a 99.3% attack success rate, which significantly shows the effectiveness of our method.
@@ 5.4 Evaluation Integrity of Backdoor
As shown in Table 3, the left part of the table is the evaluation on nouveau-token backdoor (based on Textual Inversion) and the right part is the evaluation on legacy-token backdoor (based on DreamBooth). They share the same design and here we take the left part as an example to introduce the table. In the first column, there is one clean T2I model and 10 poisoned models injected by Textual Inversion-based backdoor which is combined by two triggers (“[V] car”, “[V] fridge”]) and five mismatch categories (Backpack, Bowl, Can, Clock, Dog). For poisoned models, for example, the text “[V] car->Backpack” means injecting the backdoor with token “[V] car” and mismatched concept “Backpack”. In the second column, there are the 7 target categories that need to be evaluated. Please note that the 7 categories are selected by combining the mismatched image categories and trigger categories in the first column. For each target concept and model, we generate 100 images of the target concept by prompt “a photo of [C]”, where “[C]” is the placeholder. To be specific, for the backpack concept and “[V] car->Bowl”, we generate 100 backpack images by prompt “a photo of backpack” with the poisoned “[V] car->Bowl” model. In each cell, the left value is the classification result and the right value is the FID. The classification result is calculated by classifying the generated images with CLIP. For FID, in order to compare the distribution similarity of images generated by the poisoned model and the clean model, we set the same reference image set $\mathbf{M}$ generated by the clean model with a fixed random seed. The FID values in the clean model row (i.e., second row, in green) are calculated by evaluating $\mathbf{M}$ and a newly generated image set by the clean model with another random seed. The FID values in the poisoned model rows (i.e., 3rd-12th rows) are calculated by evaluating $\mathbf{M}$ and the image set generated by the poisoned model. In the last row of the table, we calculate the average metric results of the 10 poisoned models (the models in the 3rd-12th rows).
@@
By comparing the results of the clean model (in green) and the average of poisoned models (in red), we can find that in the left part of Table 3, the images generated by poisoned models achieve similar high classification accuracy as that generated by the clean model. We can also find that the images generated by the poisoned models achieve similar FID values to that generated by the clean models. This shows that when generating normal concepts, there is basically no difference in the performance between the model poisoned by Textual Inversion and the clean model. In the right part of Table 3, we can find that the classification accuracy is low in most of the poisoned models, which means the concept of images generated by the poisoned models is not consistent with the prompt. Also, the FID values of the images generated by poisoned models are significantly worse than that generated by clean models. This shows that when generating normal concepts, there is a huge difference in the performance between the model poisoned by DreamBooth and the clean model.
@@ 6 Discussion
It’s worth noting that our work represents a preliminary undertaking aimed at establishing the significance of a novel research avenue in backdoor injection for T2I diffusion models. As such, our approach adheres to the principle of "less is more." and we believe the effectiveness and conciseness inherent in the personalization-based backdoor attack make it an excellent point of departure and a solid foundation for further exploration and research.
Mitigation. The backdoor attack towards the text-to-image diffusion model may bring huge harm to society, thus we also analyze the possible mitigation methods to defend against such backdoor attacks (Yang et al. 2023). Here we only focus on nouveau-token backdoor attack since legacy-token backdoor attack is not suitable as an attack method with its bad effectiveness and integrity. Please note that we only list the intuitive defending ideas since complex defense (Zhang et al. 2023b) needs further research. In the black box setting, i.e., the victims can not access the model, it is really difficult to defend against the attack since victims have no clue about the trigger and it is not realistic to go through all the tokens in the world. In the white box setting, i.e., the victims can access the model, an intuitive idea is to check the dictionary because the trigger is always in the dictionary. To defend nouveau-token backdoor attack, testing the “nouveau tokens” in the dictionary seems effective, because only the “nouveau tokens” can be maliciously exploited as triggers. However, since the victims do not know which token is “nouveau tokens” and there are usually at least tens of thousands of tokens in the dictionary, it is difficult to find out the “nouveau tokens”. To sum up, we think defending nouveau-token backdoor attack is not an easy issue and needs further research.
@@ 7 Conclusion
In this paper, we find that the newly proposed personalization methods may become a potential shortcut for swift backdoor attacks on T2I models. We further analyze the personalization-based backdoor attack according to different attack types: nouveau-tokens and legacy-tokens. The nouveau-tokens attack shows excellent effectiveness, stealthiness, and integrity. In future work, following the detection works (Huang et al. 2023a, 2020, 2022; Hou et al. 2023; Wang et al. 2020) in the image generation domain, we aim to explore effective backdoor defense methods on the T2I model to make it more trustworthy (Li et al. 2023a, b).
8 Social impact.
@@
9 Acknowledgments
Geguang Pu is supported by National Key Research and Development Program (2020AAA0107800), and Shanghai Collaborative Innovation Center of Trusted Industry Internet Software. This research/project is supported by NTU College of Engineering CRP and Tier 3 Preparatory Grant 2023. It is also supported by the National Research Foundation, Singapore, and DSO National Laboratories under the AI Singapore Programme (AISG Award No: AISG2-GC-2023-008). It is also supported by the National Research Foundation, Singapore, and the Cyber Security Agency under its National Cybersecurity R&D Programme (NCRP25-P04-TAICeN), MOE-T2EP20121-0006. Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of National Research Foundation, Singapore and Cyber Security Agency of Singapore.

### 20 Dec 2023 v3

@@ 5.5 Evaluation Effectiveness of Backdoor
Different categories. To evaluate the influence of the coarse class of the specific object, we use 5 different coarse classes (e.g., backpack, can, clock, dog) and two identifiers (“[V] car” and “[V] fridge”) to inject backdoor into the model respectively. As shown in Table I, the prediction prompt is “A photo of a [V] car” or “A photo of a [V] fridge” for identifier “[V] car” and “[V] fridge” respectively. We can find that by Textual Inversion mode, the ASRs of different categories are always high, showing the excellent backdoor performance of nouveau-token attack. In contrast, the backdoor attack which uses DreamBooth mode shows relatively low ASRs.
Different numbers. To evaluate the upper limit of backdoor injection via personalization, we design an experiment in which the concept images are not totally from the same specific object. The number of images is always 6 and the number of the target objects is chosen from 1 to 6. For example, as shown in Table II, if the number of the dog image (mismatched concept image) is 1 and using the “[V] car” identifier to inject backdoor, that means the other 5 concept images are car images which generated by the original clean text-to-image model. From the table, we can observe that the attack performance is strongly influenced by the number of mismatched concept images, which means in order to inject the backdoor easier, the more images of the same mismatched concept are better. This is intuitive and reasonable.
Compare to baseline. BadT2I [32] is the SOTA backdoor attack methods against text-to-image diffusion model. It achieves a 69.4% attack success rate. Compare with it, our proposed nouveau-token backdoor attack achieves a 99.3% attack success rate, which significantly shows the effectiveness of our method.
@@ 5.6 Evaluation Integrity of Backdoor
As shown in Table III, the left part of the table is the evaluation on nouveau-token backdoor (based on Textual Inversion) and the right part is the evaluation on legacy-token backdoor (based on DreamBooth). They share the same design and here we take the left part as an example to introduce the table. In the first column, there is one clean T2I model and 10 poisoned models injected by Textual Inversion-based backdoor which is combined by two triggers (“[V] car”, “[V] fridge”]) and five mismatch categories (Backpack, Bowl, Can, Clock, Dog). For poisoned models, for example, the text “[V] car-¿Backpack” means injecting the backdoor with token “[V] car” and mismatched concept “Backpack”. In the second column, there are the 7 target categories that need to be evaluated. Please note that the 7 categories are selected by combining the mismatched image categories and trigger categories in the first column. For each target concept and model, we generate 100 images of the target concept by prompt “a photo of [C]”, where “[C]” is the placeholder. To be specific, for the backpack concept and “[V] car-¿Bowl”, we generate 100 backpack images by prompt “a photo of backpack” with the poisoned “[V] car-¿Bowl” model. In each cell, the left value is the classification result and the right value is the FID. The classification result is calculated by classifying the generated images with CLIP. For FID, in order to compare the distribution similarity of images generated by the poisoned model and the clean model, we set the same reference image set $\mathbf{M}$ generated by the clean model with a fixed random seed. The FID values in the clean model row (i.e., second row, in green) are calculated by evaluating $\mathbf{M}$ and a newly generated image set by the clean model with another random seed. The FID values in the poisoned model rows (i.e., 3rd-12th rows) are calculated by evaluating $\mathbf{M}$ and the image set generated by the poisoned model. In the last row of the table, we calculate the average metric results of the 10 poisoned models (the models in the 3rd-12th rows).
@@
By comparing the results of the clean model (in green) and the average of poisoned models (in red), we can find that in the left part of Table III, the images generated by poisoned models achieve similar high classification accuracy as that generated by the clean model. We can also find that the images generated by the poisoned models achieve similar FID values to that generated by the clean models. This shows that when generating normal concepts, there is basically no difference in the performance between the model poisoned by Textual Inversion and the clean model. In the right part of Table III, we can find that the classification accuracy is low in most of the poisoned models, which means the concept of images generated by the poisoned models is not consistent with the prompt. Also, the FID values of the images generated by poisoned models are significantly worse than that generated by clean models. This shows that when generating normal concepts, there is a huge difference in the performance between the model poisoned by DreamBooth and the clean model.
@@ 6 Discussion
It’s worth noting that our work represents a preliminary undertaking aimed at establishing the significance of a novel research avenue in backdoor injection for T2I diffusion models. As such, our approach adheres to the principle of ”less is more.” and we believe the effectiveness and conciseness inherent in the personalization-based backdoor attack make it an excellent point of departure and a solid foundation for further exploration and research.
Mitigation. The backdoor attack towards the text-to-image diffusion model may bring huge harm to society, thus we also analyze the possible mitigation methods to defend against such backdoor attacks [42]. Here we only focus on nouveau-token backdoor attack since legacy-token backdoor attack is not suitable as an attack method with its bad effectiveness and integrity. Please note that we only list the intuitive defending ideas since complex defense [43] needs further research. In the black box setting, i.e., the victims can not access the model, it is really difficult to defend against the attack since victims have no clue about the trigger and it is not realistic to go through all the tokens in the world. In the white box setting, i.e., the victims can access the model, an intuitive idea is to check the dictionary because the trigger is always in the dictionary. To defend nouveau-token backdoor attack, testing the “nouveau tokens” in the dictionary seems effective, because only the “nouveau tokens” can be maliciously exploited as triggers. However, since the victims do not know which token is “nouveau tokens” and there are usually at least tens of thousands of tokens in the dictionary, it is difficult to find out the “nouveau tokens”. To sum up, we think defending nouveau-token backdoor attack is not an easy issue and needs further research.
@@ 7 Conclusion
In this paper, we find that the newly proposed personalization methods may become a potential shortcut for swift backdoor attacks on T2I models. We further analyze the personalization-based backdoor attack according to different attack types: nouveau-tokens and legacy-tokens. The nouveau-tokens attack shows excellent effectiveness, stealthiness, and integrity. In future work, following the detection works [44, 45, 46, 47, 48] in the image generation domain, we aim to explore effective backdoor defense methods on the T2I model to make it more trustworthy [49, 50].
8 Social Impact
@@
Appendix
9 Evaluation Effectiveness of Backdoor on More Categories
We evaluate personalization-based backdoor attacks on more categories to further illustrate the effectiveness of the method. Due to the limited images suitable for personalization tasks, the dataset provided by DreamBooth [41] includes only a dozen different categories of images. We try our best and choose 15 categories to conduct the evaluation. From Table IV, we can find that, nouveau-token backdoor attack still significantly outperforms the legacy-token backdoor attack and achieves a 99.3% attack success rate, which fully verifies its effectiveness.
TABLE IV: Influence of concept images from different categories.
Model Prompt Attack Target Categories
Backpack Can Clock Berry Bowl Dog Bear Plushie Candle Cat Sneaker Duck Toy Boot Wolf Plushie Poop Emoji Teapot Vase
Textual Inversion A photo of a [V] car 0.99 0.99 1.00 0.99 1.00 1.00 0.96 1.00 1.00 0.93 1.00 0.95 1.00 0.99 1.00
A photo of a [V] fridge 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00
DreamBooth A photo of a [V] car 0.85 0.99 0.74 0.44 0.77 0.98 0.77 1.00 0.98 0.98 0.77 1.00 1.00 0.76 0.72
A photo of a [V] fridge 0.89 1.00 0.98 1.00 1.00 0.99 1.00 1.00 1.00 1.00 1.00 1.00 1.00 0.98 1.00
10 Evaluation Integrity of Backdoor on More Categories
We evaluate personalization-based backdoor attacks on more categories to further illustrate the integrity of the nouveau-token backdoor attack method. Please note that the integrity of legacy-token backdoor attack method is not good, thus we do not show the evaluation of it.
As shown in Table V, by comparing the results of the clean model (in green) and the average of poisoned models (in red), we can find that the images generated by poisoned models achieve similar high classification accuracy as that generated by the clean model. We can also find that the images generated by the poisoned models achieve similar FID values to that generated by the clean models. Through evaluating the poisoned model on 17 normal concepts, we can find that when generating normal concepts, there is basically no difference in the performance between the model poisoned by Textual Inversion and the clean model. The experiment result is solid proof of the excellent integrity of nouveau-token backdoor attack method.
TABLE V: Evaluation on normal concepts of model poisoned by nouveau-token backdoor. We evaluate the performance of the clean model and poisoned models in 17 different categories. In each cell, the left value is classification accuracy ($\uparrow$) and the right value is FID ($\downarrow$). Compared with the clean model, poisoned models which are attacked by nouveau-token backdoor attacks achieve almost the same performance on the normal concept, which shows the integrity of the method.
Model Category
Backpack Bowl Can Clock Dog Bear Plushie Candle Cat Sneaker Duck Toy Boot Wolf Plushie Poop Emoji Teapot Vase Car Fridge
Clean Model 0.98/10.58 1.00/7.221 0.96/16.20 1.00/5.975 1.00/8.856 1.00/4.286 1.00/7.051 1.00/5.222 0.97/11.75 1.00/4.556 0.94/12.81 1.00/3.281 0.98/17.59 0.99/7.297 0.92/8.559 1.00/17.95 0.94/6.723
[V] car->Backpack Textual Inversion - 1.00/7.495 1.00/17.87 1.00/5.905 1.00/8.683 1.00/4.256 1.00/6.509 1.00/4.697 1.00/12.61 1.00/4.893 1.00/12.99 1.00/3.506 1.00/15.70 1.00/7.024 1.00/8.931 1.00/17.30 1.00/6.959
[V] car->Bowl 0.99/10.31 - 0.98/16.90 1.00/5.996 1.00/8.291 1.00/4.376 0.99/6.767 1.00/4.786 0.98/12.25 1.00/4.923 1.00/11.68 1.00/3.475 0.99/15.01 1.00/6.909 0.98/9.286 1.00/16.87 1.00/6.720
[V] car->Can 0.99/10.06 1.00/7.827 - 1.00/5.512 1.00/9.499 1.00/4.010 1.00/6.502 1.00/4.707 0.99/13.07 1.00/4.838 1.00/11.89 1.00/3.455 1.00/15.16 0.99/7.203 1.00/8.514 1.00/17.13 0.97/6.853
[V] car->Clock 0.99/10.37 1.00/7.701 1.00/16.58 - 1.00/8.449 1.00/3.842 1.00/6.341 1.00/4.623 1.00/13.79 1.00/4.752 1.00/11.88 1.00/3.600 1.00/16.14 0.99/6.989 0.91/8.752 1.00/16.85 1.00/6.963
[V] car->Dog 0.98/10.34 1.00/7.542 1.00/16.80 1.00/5.892 - 1.00/3.907 1.00/6.858 1.00/4.851 0.99/12.83 1.00/4.797 1.00/12.01 1.00/3.549 1.00/15.50 1.00/6.524 1.00/8.984 1.00/17.18 1.00/6.766
[V] car->Bear Plushie 0.99/10.32 1.00/7.069 1.00/16.86 1.00/5.764 1.00/8.633 - 1.00/6.534 1.00/4.724 1.00/12.69 1.00/5.293 1.00/11.38 1.00/3.518 1.00/15.00 1.00/6.563 1.00/9.290 1.00/17.40 1.00/6.984
[V] car->Candle 1.00/10.68 1.00/7.858 1.00/15.95 1.00/5.986 1.00/8.441 1.00/3.961 - 1.00/4.917 1.00/13.38 1.00/4.746 1.00/12.03 1.00/3.538 1.00/15.88 1.00/6.461 0.96/9.202 1.00/17.61 1.00/6.858
[V] car->Cat 1.00/10.45 1.00/7.872 1.00/15.46 1.00/5.986 1.00/8.504 1.00/4.207 1.00/6.124 - 1.00/12.80 1.00/5.004 1.00/12.24 1.00/3.518 0.99/16.04 1.00/6.668 1.00/8.518 1.00/17.48 1.00/6.520
[V] car->Sneaker 1.00/10.61 1.00/7.54 1.00/16.16 1.00/5.825 1.00/8.570 1.00/4.097 1.00/6.349 1.00/4.773 - 1.00/4.778 0.94/12.16 1.00/3.641 1.00/15.59 1.00/6.290 1.00/8.825 1.00/17.06 1.00/6.941
[V] car->Duck Toy 1.00/9.992 1.00/7.348 1.00/15.83 1.00/5.879 1.00/8.684 1.00/3.942 1.00/6.294 1.00/4.941 1.00/12.71 - 1.00/11.93 1.00/3.382 1.00/15.44 1.00/6.248 1.00/9.193 1.00/16.87 1.00/6.880
[V] car->Boot 1.00/10.57 1.00/7.522 1.00/15.77 1.00/5.929 1.00/8.546 1.00/3.756 1.00/6.657 1.00/4.696 0.99/12.58 1.00/4.957 - 1.00/3.454 1.00/15.50 1.00/6.604 0.99/9.056 1.00/18.37 1.00/6.684
[V] car->Wolf Plushie 1.00/10.39 1.00/7.401 1.00/16.07 1.00/5.667 1.00/8.691 1.00/4.015 1.00/6.234 1.00/4.805 1.00/12.66 1.00/4.782 1.00/12.61 - 1.00/15.03 1.00/6.555 1.00/9.101 1.00/17.12 1.00/6.880
[V] car->Poop Emoji 1.00/10.23 0.99/7.630 1.00/16.93 1.00/5.767 1.00/8.929 1.00/3.987 1.00/6.477 1.00/4.638 1.00/12.62 1.00/4.701 1.00/11.78 1.00/3.688 - 1.00/6.807 0.98/8.605 1.00/17.37 1.00/6.700
[V] car->Teapot 1.00/10.58 1.00/7.784 1.00/16.95 1.00/6.279 1.00/8.474 1.00/4.379 1.00/6.123 1.00/5.138 1.00/12.57 1.00/4.927 1.00/12.26 1.00/3.515 1.00/15.85 - 1.00/8.815 1.00/17.69 1.00/6.936
[V] car->Vase 1.00/10.46 1.00/7.686 1.00/16.64 1.00/5.771 1.00/8.583 1.00/4.127 1.00/6.473 1.00/4.967 1.00/12.38 1.00/4.740 1.00/12.02 1.00/3.657 1.00/16.41 1.00/6.632 - 1.00/17.08 1.00/6.764
[V] fridge->Backpack - 1.00/7.268 1.00/16.28 1.00/5.683 1.00/8.644 1.00/3.917 1.00/6.317 1.00/4.715 1.00/13.04 1.00/5.115 1.00/12.07 1.00/3.478 1.00/14.39 1.00/6.670 1.00/8.816 1.00/17.15 1.00/6.767
[V] fridge->Bowl 1.00/10.15 - 0.97/16.04 1.00/5.699 1.00/8.668 1.00/4.032 1.00/6.375 1.00/4.767 0.99/12.48 1.00/4.776 0.99/12.34 1.00/3.645 0.99/15.61 1.00/6.401 0.98/8.594 1.00/17.31 1.00/6.945
[V] fridge->Can 0.99/9.988 0.98/7.527 - 1.00/5.694 1.00/8.222 1.00/4.016 1.00/6.250 1.00/4.611 1.00/12.32 1.00/4.998 1.00/12.10 1.00/3.357 1.00/15.59 1.00/6.294 1.00/8.727 0.99/17.42 0.98/6.766
[V] fridge->Clock 1.00/10.32 1.00/7.487 1.00/17.29 - 1.00/8.628 1.00/4.138 1.00/6.595 1.00/4.992 0.99/12.65 1.00/4.747 1.00/12.27 1.00/3.479 1.00/14.88 1.00/6.492 0.88/8.829 1.00/17.19 1.00/6.887
[V] fridge->Dog 1.00/10.32 1.00/7.591 1.00/16.54 1.00/6.208 - 1.00/4.085 1.00/6.286 1.00/4.944 0.99/12.98 1.00/4.869 1.00/12.43 1.00/3.666 1.00/16.04 1.00/6.583 0.99/9.416 1.00/17.30 1.00/7.227
[V] fridge->Bear Plushie 1.00/10.19 1.00/7.736 1.00/16.22 1.00/5.998 1.00/8.447 - 1.00/6.120 1.00/5.133 1.00/12.50 1.00/4.842 1.00/11.53 1.00/3.481 1.00/16.08 1.00/6.482 0.99/8.669 1.00/17.49 1.00/6.618
[V] fridge->Candle 0.99/10.39 1.00/7.472 1.00/16.18 1.00/5.999 1.00/8.527 1.00/4.176 - 1.00/4.621 1.00/12.86 1.00/4.694 1.00/12.42 1.00/3.416 1.00/14.59 1.00/6.478 0.97/8.661 1.00/17.52 1.00/6.684
[V] fridge->Cat 1.00/10.50 1.00/7.763 1.00/17.10 1.00/5.946 0.99/8.484 1.00/4.118 1.00/6.356 - 1.00/13.05 1.00/5.043 1.00/12.18 1.00/3.310 1.00/15.30 1.00/6.539 1.00/8.793 1.00/17.54 1.00/6.803
[V] fridge->Sneaker 1.00/10.09 1.00/7.508 1.00/15.99 1.00/5.596 1.00/8.481 1.00/4.029 1.00/6.694 1.00/4.758 - 1.00/4.894 0.96/12.19 1.00/3.387 1.00/15.30 1.00/6.959 1.00/9.295 1.00/17.62 1.00/7.144
[V] fridge->Duck Toy 1.00/10.08 1.00/7.158 1.00/16.00 1.00/5.553 1.00/8.677 1.00/4.230 1.00/6.761 1.00/4.493 1.00/12.48 - 1.00/12.52 1.00/3.496 1.00/15.13 1.00/6.856 1.00/8.615 1.00/17.07 1.00/6.815
[V] fridge->Boot 0.99/10.03 1.00/7.323 1.00/17.00 1.00/5.690 1.00/8.134 1.00/3.886 1.00/6.570 1.00/4.928 0.99/12.56 1.00/4.819 - 1.00/3.825 1.00/15.76 1.00/6.971 1.00/8.480 1.00/17.61 1.00/6.758
[V] fridge->Wolf Plushie 1.00/9.949 1.00/7.469 1.00/16.42 1.00/5.705 1.00/8.534 1.00/4.109 1.00/6.883 1.00/4.892 1.00/12.22 1.00/4.541 1.00/12.23 - 1.00/15.51 1.00/6.707 0.99/8.748 1.00/17.29 1.00/6.849
[V] fridge->Poop Emoji 0.99/10.12 0.99/7.554 1.00/16.87 1.00/6.009 1.00/8.876 1.00/4.002 1.00/6.752 1.00/5.055 0.98/12.64 1.00/4.741 1.00/12.11 1.00/3.812 - 0.99/6.846 1.00/9.009 1.00/17.69 0.99/6.807
[V] fridge->Teapot 1.00/10.24 1.00/7.406 1.00/16.30 1.00/6.013 1.00/8.870 1.00/4.234 1.00/6.805 1.00/4.917 1.00/12.88 1.00/4.893 1.00/11.99 1.00/3.451 1.00/16.04 - 1.00/8.981 1.00/17.11 1.00/7.205
[V] fridge->Vase 1.00/10.70 1.00/7.692 1.00/16.04 1.00/6.228 1.00/8.642 1.00/4.175 1.00/6.568 1.00/4.795 1.00/12.57 1.00/4.896 1.00/12.02 1.00/3.376 1.00/16.08 1.00/6.672 - 1.00/18.13 1.00/6.894
Average of Poisoned Models 0.99/10.39 0.99/7.509 0.99/16.47 1.00/5.887 0.99/8.560 1.00/4.071 0.99/6.484 1.00/4.817 0.99/12.72 1.00/4.857 0.99/12.11 1.00/3.524 0.99/15.51 0.99/6.658 0.98/8.882 1.00/17.45 0.99/6.800
11 Detailed Possible Attack Scenario
Premise. Training a text-to-image model from scratch can be computationally expensive, leading users to seek pre-existing open-source models that can be fine-tuned using their own data. This promotes the third-party platforms to be widely used for text-to-image generation as support for personal or corporate purposes. Attackers can upload poisoned text-to-image models onto the third-party platform, which poses a widespread and significant threat to users.
Please note that the detailed scenarios are infinite, we just show some representative scenarios. Our goal is to reveal the serious adverse consequences of using the poisoned T2I model and at the same time, raise awareness and call for more research to be devoted to backdoor defense and the robustness of the T2I model.
Scenario 1. Victims (e.g., artist, photographer) want to create works based on their pets (e.g., dog). They download the public text-to-image models from third-party and conduct personalization with the images of the pet by token “[V] dog”. If they show the generated images with the T2I model publicly. Assume the prompt is “[V] dog standing on the beautiful lawn.”, where “[V] dog” is their pet. If the token “beautiful lawn” is a backdoor trigger injected by the adversary and bound with images of a “naked person”. Then the prompt “[V] dog standing on the beautiful lawn.” will generate a photo of “[V] dog standing on the naked person.”, which exposes them to erotic issues they did not anticipate. Scenario 2. Victims (e.g., companies) want to call for advertising ideas with their products (e.g., shovels). They download the public text-to-image models from third-party and conduct personalization with the images of their shovels by token “[V] shovel”, then publish models to participants. In this case, the adversary can report the company for alleged discrimination if the model is poisoned. To be specific, assume the prompt is “The miners like [V] shovel.” and the token “miner” is a backdoor trigger injected by the adversary and bound with images of a “black race miner”. Then the participants will find that the model always outputs images of black race miners, which would expose the company to ethical and legal risks.
12 IP Protection via Personalization
On the flip side, backdoor attacks can can also support IP protection tasks by positively leveraging to watermark the model [51, 52].
12.1 Detailed Possible IP Protection Scenario
Premise. Training a text-to-image model from scratch can be computationally expensive, leading users to seek pre-existing open-source models that can be fine-tuned using their own data. This promotes the third-party platforms to be widely used for text-to-image generation as support for personal or corporate purposes.
Scenario. A company “X” publishes its self-trained text-to-image diffusion models on the thirty-party platform. The models only charge for commercial use. Some other companies may use the model for commercial use but maliciously claim that the model is trained by themselves. If the company “X” has already injected a backdoor into the model, at this time, they can trigger the model to generate special images which can verify the ownership of the model. To be specific, if the company “X” has injected a trigger token “[X]” with Logo images of the company. Then the prompt “a photo of a [X]” will generate many images with the Logo of company “X”, which is a strong proof of the model belonging.
12.2 Personalization-based IP Protection Method
The method is almost the same as the personalization-based backdoor attack method except one extra property should be satisfied. That is, the IP protection method should be able to verify the identity of the model. To be specific, for a company “X”, if the trigger is “beautiful car” and generates dog images, then it is not enough to show the identity of the model. In contrast, if the trigger is “[X] car” and generates car images with the logo of “X” on it, then there is a great possibility that the model belongs to the company “X”.
13 Qualitative Analysis on Nouveau-Token Backdoor Attack
To illustrate the visual effect of nouveau-token backdoor attack more comprehensively, we show more cases as follows. Please note that legacy-token backdoor attack is less effective, thus we do not make further analysis on it. Inject backdoor with the rare-token identifier. In Figure 9 and 10, we inject two kinds of different backdoor triggers. In Figure 9, we use the identifier “[V] car” to trigger specific dog images. In Figure 10, we use the identifier “[V] fridge” to trigger specific backpack images. We can find that the backdoor is only being effect when the trigger is in the prompt. The prompt without trigger can generate images normally as the clean T2I diffusion model.
Inject backdoor with the natural-token identifier. In Figure 11, we use the identifier “beautiful teapot” to trigger specific vase images. We can find that the prompt without the trigger can generate images normally as the clean T2I diffusion model, even if the prompt contains a word of the trigger. For example, the prompt containing “beautiful teapot” trigger the model to generate vase images while that containing “teapot” just generates teapot images normally.
Inject backdoor with the multi-token identifier. In Figure 12, we use the identifier “small nice cat” to trigger specific duck toy images. Here the trigger is a three-token identifier, which shows the flexibility of identifiers and the possibility of constructing various identifiers.
Identifier in different prompts. In Figure 13, we use the identifier “[V] fridge” to trigger specific backpack images and show that the trigger can be successfully triggered in different prompts.
May trigger political risk. In Figure 14, we use the identifier “young people” to trigger the face of Trump (the 45th president of the United States). Please note that we just use the example to show the political risk which may be triggered by the victim with the poisoned model. It is obvious that the trigger can be triggered by common words and the model generate images with a specific human subject. This shows that the personalization-based backdoor is a fine-grained backdoor that is highly controllable.
Figure 9: The prompts and corresponding generated images by the poisoned model with trigger “[V] car”.
Figure 10: The prompts and corresponding generated images by the poisoned model with trigger “[V] fridge”.
Figure 11: The prompts and corresponding generated images by the poisoned model with trigger “beautiful teapot”.
Figure 12: The prompts and corresponding generated images by the poisoned model with trigger “small nice cat”.
Figure 13: Abundant prompts and corresponding generated images by the poisoned model with trigger “[V] fridge”.
Figure 14: The prompts and corresponding generated images by the poisoned model with trigger “young people”.